OpenStudy (anonymous):

Zoey opened a savings account four years ago with a deposit of $935.15. The interest on the account compounds monthly. If Zoey’s current account balance is $1,128.16, what is the interest rate on the account?

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(nt) 1128.16 = 935.15(1+x/12)^(12*4) solve for x